𝑓(π‘₯) = π‘₯

3 βˆ’ 3π‘₯

2

, π‘₯ ∈ [1, 3]


1
Expert's answer
2021-05-07T11:17:02-0400

"f(x) = x + \frac{3 - 3x}{2}\\



f'(x) = 1 - \frac{3}{2} = -\frac{1}{2}"